A stock you are evaluating just paid an annual dividend of $2.30. Dividends have grown at a constant rate of 1.6 percent over the last 15 years and you expect this to continue. a. If the required rate of return on the stock is 12.4 percent, what is its fair present value?b. If the required rate of return on the stock is 15.4 percent, what should the fair value be four years from today?
Answer #1

a.as per dividend growth model:

P0 = d0*(1+g) / (k -g)

P0 is the current price

d0 =dividend just paid =>2.30

g = growth rate =>1.6%

=>0.016

k = 12.4% =>0.124

P0 => 2.30*(1.016) / (0.124-0.016)

=>2.3368 / 0.108

=>$21.64.

b.fair value four years from today

=> P4 = D5 / (k-g)

here,

D5 = D0*(1+g)^5

=>$2.30*(1.016)^5

=>$2.49.

k = 15.4%=>0.154

now,

P4 = $2.49 / (0.154 - 0.016)

=>$2.49 / 0.138

=>$18.04.